OpenStudy (anonymous):

(xy^9 / 3y^-2) • (-7y / 21x^5)

OpenStudy (anonymous):

\[\frac{ xy ^{9} }{ 3y ^{-2} } \times \frac{ -7y }{ 21x ^{5} }\]

OpenStudy (ivettef365):

first you can simplify, -7/21 = ??? to simplify variables with exponents on division it is (x^(a-b)) so you have y^9 and y^-2 so y^(9-(-2)) = y^11

OpenStudy (anonymous):

Yes. I already have an answer, but I was hoping that if someone would be willing to work through the problem, we would get the same answer. First thing I did was simplify -7/21 to -1/3, then got y^11. I ended up with \[\frac{ xy ^{11} }{ 3 } \times \frac{ -y }{ 3x ^{5} }\] Correct?

OpenStudy (ivettef365):

ok, now you can simplify the x from the top with the x from the bottom and you have y^11 -y ------- x ----- 3 3x^4

OpenStudy (ivettef365):

now multiply and what do you get ?

jhonyy9 (jhonyy9):

so do you understand it ?

jhonyy9 (jhonyy9):

courage

OpenStudy (anonymous):

\[\frac{ -2y ^{12} }{ 9x ^{4} }\]

OpenStudy (ivettef365):

actuallry (1) x (-1) = -1

OpenStudy (ivettef365):

so you have -y^12 ------- 9x^4

OpenStudy (anonymous):

Yeah, I had -y on my paper. I don't know why I decided to include a 2.

OpenStudy (anonymous):

Yeah, that's what I had the first time too. I just wasn't positive, so I needed some reassurance. Thanks for walking me through it!

OpenStudy (ivettef365):

yw :)